# Token Is All You Need: Cognitive Planning through Sparse Intent Alignment

### v1

@@ IV-B Main Results
Table IV-B summarizes our ablation results on nuPlan, along with representative state-of-the-art (SOTA) methods on nuScenes. TIWM achieves strong performance through task-driven semantic alignment, without dense reconstruction or auxiliary supervision.
Table 1: Ablation results on nuPlan, with SOTA methods on nuScenes shown for reference. Our framework achieves high accuracy via sparse intent alignment, without explicit reconstruction. Numerical comparisons across datasets are indicative only. Method Future Intent Self-Supervision Type ADE (m) FDE (m) Params (M) Methods evaluated on nuScenes (reference only) UniAD [1] $\times$ Multi-task Auxiliary 1.03 1.65 $\sim$80 VAD-Base [2] $\times$ Multi-task Auxiliary 1.22 1.98 $\sim$55 SSR [4] $\times$ Dense Reconstruction 0.75 1.36 $\sim$32 Our ablation study on nuPlan (main results) Current token, no intent loss $\times$ None 0.548 0.993 14.2 Current token, with intent loss $\times$ Sparse Intent Reconstruction 0.554 1.007 14.2 Future token, no intent loss Task-driven Semantic Alignment 0.479 0.864 14.2 Future token, with intent loss Sparse Intent Reconstruction 0.489 0.890 14.2
Note: Parameter counts for SSR, UniAD, and VAD are estimated since exact values are not reported. SSR uses dense BEV reconstruction (“we supervise directly using an L2 loss with the real future BEV feature” [4]), whereas TIWM aligns only 16 sparse intent tokens—conceptually distinct from feature-level reconstruction. Within nuPlan, the Future token, no intent loss setting attains 0.479 m ADE—13.5% better than its reconstruction-based variant with current token.
@@
(1) Minimal complexity yields strong performance. Even without future prediction, TIWM achieves 0.548 m ADE, comparable to or surpassing prior benchmarks on nuScenes. This demonstrates that semantically reliable inputs enable high-quality planning without dense reconstruction. Within the same setup, removing future prediction increases ADE by only 0.006 m, showing that additional complexity contributes little.
(2) Future-conditioned decoding enhances reasoning. Conditioning trajectory generation on predicted future tokens—rather than reconstructed BEV features—reduces ADE to 0.479 m, a 12.6% gain over the baseline. This validates that planning via semantic imagination, rather than pixel-level reconstruction, drives performance.
(3) Explicit reconstruction loss hinders learning. In contrast to SSR’s reconstruction-based self-supervision, adding an intent-alignment term slightly degrades performance (0.489 m vs. 0.479 m ADE) and slows convergence (best epoch 43 vs. 40). Trajectory supervision alone provides a cleaner and more effective optimization signal.
@@
The key distinction between TIWM and SSR lies in their self-supervision philosophy. SSR [4] enforces pixel-level alignment by reconstructing future BEV features under auxiliary trajectory guidance, effectively coupling perception, prediction, and planning through dense supervision. In contrast, TIWM adopts a task-driven, token-level semantic alignment strategy—reconstructing only the compact intent representation that is directly optimized for the planning objective. This design decouples low-level feature reconstruction from high-level decision semantics, resulting in a more efficient and cognitively interpretable learning process.
@@ IV-C Ablation Study
  • Current tokens, no intent loss: decode from current tokens only, without future prediction or reconstruction.
  • Current tokens, with intent loss: decode from current tokens with auxiliary reconstruction supervision.
  • Future tokens, no intent loss: decode from current and predicted future tokens without reconstruction.
  • Future tokens, with intent loss: decode from current and predicted future tokens with explicit L2 reconstruction.
Fig. 2 reports validation ADE and training loss across configurations. All models improve steadily up to epoch 50, with the best performance typically around epochs 40–45, suggesting that longer training enhances temporal reasoning.
Figure 2: Training curves for four configurations: validation ADE (left, lower is better) and training loss (right). “No intent loss” variants consistently outperform their reconstruction-based counterparts after epoch 20.
@@
IV-D Discussion
Our experiments reveal three core principles of cognitive planning:
@@ IV-D Discussion
(1) Task-driven semantic alignment surpasses pixel-level reconstruction. The consistent superiority of the No Intent Loss configuration demonstrates that semantic alignment driven solely by the trajectory objective outperforms explicit reconstruction. This aligns with human cognition, where attention selectively prioritizes task-relevant cues rather than exhaustively reconstructing the visual scene. As Li et al. [4] noted, reconstruction-heavy paradigms often require multi-objective optimization across perception, prediction, and planning, leading to unstable convergence. Similarly, Zheng et al. [10] observed that overly entangled objectives hinder error recovery and safety generalization in autonomous driving. By focusing learning on decision-critical semantics, TIWM achieves greater stability, efficiency, and interpretability.
@@
(2) “Temporal fuzziness” is a feature, not a flaw. The model learns what to focus on rather than when, maintaining probabilistic attention over key semantic elements. This flexibility enables robust reasoning under uncertainty and avoids overfitting to strict temporal alignment—an inherent limitation of frame-based world models.
@@
(3) Planning through imagination outperforms reactive policies. The 12.6% performance gain from future-conditioned decoding confirms that intelligence arises from anticipating multiple futures before acting. This addresses the Markovian limitation of VLA systems [12], demonstrating that intent evolution provides a computational form of mental simulation that supports foresightful planning.
@@
Together, these insights mark a shift from reconstructing the world to understanding the world. Instead of expending computation on dense reconstruction—as in SSR [19]—TIWM achieves stronger results through lightweight, task-driven semantic alignment. This bridges two long-divided paradigms in embodied AI: world models emphasizing exhaustive scene synthesis and VLA systems limited to instantaneous decision-making. We show that belief–intent co-evolution, not pixel fidelity, underpins effective planning.
The implications of this work extend beyond autonomous driving. The “token is all you need” principle offers a general cognitive lens for embodied intelligence. In robotic manipulation, sparse intent tokens could represent affordances rather than full object geometry; in conversational AI, they could encode high-level discourse states instead of exhaustive linguistic context—focusing computation on the semantics that truly drive behavior.
More broadly, these findings suggest that intelligence does not arise from reconstructing every detail of the environment, but from learning to capture and reason over the information that matters most for the task at hand. TIWM provides not only an efficient planning framework but also an initial step toward unifying representation, intent, and prediction—a direction consistent with cognitive theories of internal world modeling. By grounding decision-making in tokenized, semantically aligned representations, TIWM bridges the gap between data-driven perception and cognitively inspired reasoning, pointing toward a new generation of agents that plan through understanding rather than reaction.
V Conclusion
@@
We challenge the prevailing assumption that high-fidelity future scene reconstruction is necessary for effective end-to-end planning. Through extensive experiments on the nuPlan benchmark (720 scenarios, over 11,000 samples), we show that task-driven semantic alignment—rather than pixel-level reconstruction—can achieve high planning accuracy while remaining efficient and interpretable. A lightweight planner operating on only 16 discrete intent tokens reaches 0.479 m ADE, demonstrating that strong performance can be achieved without dense reconstruction or auxiliary supervision, thus validating the effectiveness of our minimal, cognitively inspired design.
@@
This work establishes a new paradigm in autonomous driving: planning through understanding. We demonstrate that:
1) High-quality inputs enable powerful yet simple planners. Reliable perception allows the planner to achieve strong performance without auxiliary reconstruction losses, suggesting that high accuracy does not require complex or over-parameterized models.
2) Future-aware decoding enables foresightful behavior. Conditioning trajectory generation on predicted intent tokens—rather than reconstructed scenes—yields a 12.6% improvement, illustrating that imagination-driven reasoning underlises intelligent control.
3) Explicit reconstruction loss offers no benefit. Explicit reconstruction loss offers no benefit; it constrains the model. Instead, a clean optimization signal focused solely on the final planning outcome naturally yields robust and generalizable intent representations.
The emergent property of ”temporal fuzziness”—the ability to focus on what matters semantically rather than rigidly aligning to when—further highlights the cognitive flexibility of our approach. This enables human-like adaptability under uncertainty.
Beyond autonomous driving, TIWM suggests a broader principle for embodied intelligence: true autonomy arises not from reconstructing the world in full fidelity, but from understanding its semantic essence and imagining possible futures. By unifying representation, intent, and prediction into a single, token-based framework, TIWM bridges the gap between data-driven perception and cognitively inspired reasoning. Future work will explore closed-loop evaluation, token scalability, and the generalization of belief–intent co-evolution to other domains. We view this as a step toward agents that plan through understanding, suggesting that sparse, task-aligned tokenization may represent a fundamental mechanism for efficient cognition in artificial systems.

### v2

@@ IV-B Main Results
Table I summarizes our ablation results on nuPlan, along with representative state-of-the-art (SOTA) methods on nuScenes. TIWM achieves strong performance through task-driven semantic alignment, without dense reconstruction or auxiliary supervision.
Table I: Ablation results on nuPlan, with SOTA methods on nuScenes shown for reference. Our framework achieves high accuracy via sparse intent alignment, without explicit reconstruction. Numerical comparisons across datasets are indicative only. Method Future Intent Self-Supervision Type ADE (m) FDE (m) Params (M) Methods evaluated on nuScenes (reference only) UniAD [huPlanningOrientedAutonomousDriving2023a] $\times$ Multi-task Auxiliary 1.03 1.65 $\sim$80 VAD-Base [jiangVADVectorizedScene2023a] $\times$ Multi-task Auxiliary 1.22 1.98 $\sim$55 SSR [liNavigationGuidedSparseScene2025] $\times$ Dense Reconstruction 0.75 1.36 $\sim$32 Our ablation study on nuPlan (main results) Current token without intent loss $\times$ None 0.487 0.855 14.2 Current token with intent loss $\times$ Sparse Intent Reconstruction 0.459 0.810 14.2 Future token without intent loss Task-driven Semantic Alignment 0.382 0.699 14.2 Future token with intent loss Sparse Intent Reconstruction 0.492 0.944 14.2 Note: Parameter counts for SSR, UniAD, and VAD are estimated since exact values are not reported. SSR uses dense BEV reconstruction (“we supervise directly using an L2 loss with the real future BEV feature” [liNavigationGuidedSparseScene2025]), whereas TIWM aligns only 16 sparse intent tokens—conceptually distinct from feature-level reconstruction. Within nuPlan, the Future token without intent loss setting attains 0.382 m ADE—16.8% better than its reconstruction-based variant with current token. Dataset Context: nuScenes and nuPlan share a common data origin but serve different purposes. nuPlan is specifically designed as a planning benchmark, featuring complex urban scenarios curated to stress-test long-horizon reasoning and interactive decision-making, making it more challenging than nuScenes for evaluating pure planning performance. In contrast, many two-stage E2E methods on nuPlan report closed-loop metrics that often rely on rule-based post-processing to ensure safety, obscuring their open-loop trajectory prediction capabilities. Our results on nuPlan are reported in open-loop ADE/FDE, providing a clean evaluation of the core planning module without such post-processing.
@@
(1) Minimal complexity yields strong performance. Even without future prediction, TIWM achieves 0.487 m ADE, comparable to or surpassing prior benchmarks on nuScenes. This demonstrates that semantically reliable inputs enable high-quality planning without dense reconstruction.
(2) Future-conditioned decoding enhances reasoning. Conditioning trajectory generation on predicted future tokens—rather than reconstructed BEV features—reduces ADE to 0.382 m, a 21.6% gain over the baseline. This validates that planning via semantic imagination, rather than pixel-level reconstruction, drives performance.
(3) Explicit reconstruction loss hinders learning. Incorporating future token decoding alongside an intent-alignment term slightly degrades performance (ADE: 0.492 m vs. 0.382 m) and introduces instability in validation performance. Furthermore, this configuration yields a higher training loss compared to the other three variants, suggesting a gradient conflict arising from the competing objectives of trajectory reconstruction and intent alignment.
@@
The key distinction between TIWM and SSR lies in their self-supervision philosophy. SSR [liNavigationGuidedSparseScene2025] enforces pixel-level alignment by reconstructing future BEV features under auxiliary trajectory guidance, effectively coupling perception, prediction, and planning through dense supervision. In contrast, TIWM adopts a task-driven, token-level semantic alignment strategy—reconstructing only the compact intent representation that is directly optimized for the planning objective. This design decouples low-level feature reconstruction from high-level decision semantics, resulting in a more efficient and cognitively interpretable learning process.
@@ IV-C Ablation Study
  • Current tokens without intent loss: decode from current tokens only, without future prediction or reconstruction.
  • Current tokens with intent loss: decode from current tokens with auxiliary reconstruction supervision. $\mathcal{L}_{\text{intent}}=100$;
  • Future tokens without intent loss: decode from current and predicted future tokens without reconstruction.
[…]
Fig. 2 reports validation ADE and training loss across configurations. All models improve steadily up to epoch 100, demonstrating strong potential for further convergence. However, prolonged training incurs significant computational costs. This trend suggests that longer training enhances temporal reasoning, albeit at the expense of efficiency.
Figure 2: Validation ADE (left) and training loss (right) versus epoch across four configurations. All runs converge stably, with late-epoch best ADEs between 0.382 and 0.492 m. Early-epoch behavior is configuration-dependent: with current tokens, intent conditioning yields modest early gains (lower mean ADE within the first 100 epochs), whereas with future tokens, the non-intent variant reaches lower ADE faster. The overall best minimum is obtained by the future-token model without intent.
@@
IV-D Convergence Study
As shown in Table II, the model’s performance improves with increasing training epochs, albeit with diminishing returns. The consistent downward trend in both validation ADE and training loss (Fig. 3) suggests that the model possesses a remarkable capacity for prolonged learning, continuously refining its internal representations even after 1000 epochs.
This persistent convergence is not merely a numerical phenomenon; it reflects a deeper principle of cognitive consistency. We propose that the sparse tokens within the intent chain function as high-level cognitive units, embodying a dual nature:
  • Belief: They serve as a compressed representation of the current environment, encoding task-relevant semantics.
  • Intent: They represent the agent’s imagined future goals, guiding the planning process.
We hypothesize that during training, these token units undergo a dynamic equilibrium between their ”belief” and ”intent” states—a process we metaphorically describe as a ”quantum state transition.” This continuous balancing act represents a more fundamental form of learning than conventional supervised or reinforcement learning, as it involves the self-organization of an internal world model.
This interpretation also offers a compelling explanation for the biological phenomena of neural replay [huangReplaytriggeredBrainwideActivation2024] and memory consolidation [squireMemoryConsolidation2015]: just as the brain replays past experiences to strengthen memories and refine decision-making policies, our model’s long-term training can be seen as a form of ”computational replay,” where the system repeatedly refines its internal belief-intent space to achieve greater robustness and generalization.
The convergence curve thus reveals not just a decline in error, but the emergence of an implicit, natural regularization within the system—a self-organizing process that embodies the very essence of adaptive, cognition-inspired intelligence. This is a emergence of cognitive intelligence.
Figure 3: Training dynamics for the “Future token without intent loss” configuration. Left: Validation ADE (m) converges to a minimum of 0.263 m at epoch 860. Right: Training loss decreases steadily with epoch, indicating continuous optimization. This sustained convergence demonstrates the model’s capacity for prolonged learning and its potential for further performance gains.
Table II: Performance Evolution of TIWM under Prolonged Training. Training Epochs Future Intent Self-Supervision Type ADE (m) FDE (m) Improvement vs 50ep (%) 50 Task-driven Semantic Alignment 0.431 $\pm$ 0.641 0.798 $\pm$ 1.199 - 100 Task-driven Semantic Alignment 0.382 $\pm$ 0.584 0.699 $\pm$ 1.098 11.4% 300 Task-driven Semantic Alignment 0.299 $\pm$ 0.541 0.535 $\pm$ 1.009 30.6% 500 Task-driven Semantic Alignment 0.281 $\pm$ 0.546 0.506 $\pm$ 1.017 34.8% 1000 Task-driven Semantic Alignment 0.263 $\pm$ 0.558 0.471 $\pm$ 1.013 39.0% 3000 Task-driven Semantic Alignment 0.262 $\pm$ 0.567 0.463 $\pm$ 1.022 39.2% Note: This table quantifies the model’s exceptional capacity for continuous learning over extended training. The consistent improvement from 50 to 3000 epochs on the nuPlan validation set validates our core hypothesis: that task-driven semantic alignment enables deep, persistent refinement of internal representations, leading to superior generalization. The diminishing returns after 1000 epochs suggest a natural plateau, yet the system remains far from convergence, highlighting its robustness and adaptability.
V Discussion
Our experiments reveal four foundational principles of cognitive planning that collectively challenge the dominant paradigms in end-to-end autonomous driving.
@@ V Discussion
(1) Task-driven semantic alignment surpasses pixel-level reconstruction. The consistent superiority of the Future token without intent loss configuration (0.382 m ADE) over its reconstruction-based counterpart (0.492 m ADE) demonstrates that explicit alignment with future BEV features is not only unnecessary but detrimental. This aligns with human cognition, where attention prioritizes task-relevant cues rather than exhaustively reconstructing the visual scene. As SSR [liNavigationGuidedSparseScene2025] notes, reconstruction-heavy paradigms often couple perception, prediction, and planning through dense supervision, leading to unstable convergence and entangled objectives [zhengSafeOfflineReinforcement2024]. By focusing learning solely on the final planning outcome, TIWM decouples low-level feature generation from high-level decision semantics, yielding a more stable, efficient, and interpretable system.
@@
(2) “Temporal fuzziness” is a cognitive feature, not a flaw. The model learns what to focus on rather than when, maintaining probabilistic attention over key semantic elements across time. This flexibility enables robust reasoning under uncertainty and avoids overfitting to strict temporal alignment—an inherent limitation of frame-based world models. Temporal fuzziness is not noise; it is an adaptive mechanism that mirrors the brain’s ability to reason about events in a temporally coarse manner, prioritizing semantic relevance over precise timing.
@@
(3) Planning through imagination outperforms reactive policies. The 21.6% performance gain from future-conditioned decoding validates that intelligence arises not from reacting to the present, but from anticipating multiple futures before acting. This directly addresses the Markovian limitation of VLA systems [bansalChauffeurNetLearningDrive2018], demonstrating that autoregressive intent chains provide a computational form of mental simulation [decetyPowerSimulationImagining2006, mattarPlanningBrain2022] that supports foresightful planning.
(4) Cognitive consistency learning enables persistent self-optimization. The model’s continuous performance gains over 3000 epochs—converging to 0.262 m ADE—reveal an emergent self-organizing process. We propose that sparse tokens function as high-level cognitive units, dynamically balancing belief (current perception) and intent (future imagination). This equilibrium, which we metaphorically describe as a “quantum state transition,” mirrors biological phenomena like neural replay [huangReplaytriggeredBrainwideActivation2024] and memory consolidation [squireMemoryConsolidation2015]. Long-term training thus acts as “computational replay,” where the system repeatedly refines its internal belief–intent space to achieve greater robustness—a hallmark of adaptive, cognition-inspired intelligence.
On Cross-Dataset Comparisons. While our results are reported on nuPlan and prior SOTA methods on nuScenes, the comparison is meaningful. nuPlan is explicitly designed as a planning-focused benchmark, featuring complex, interactive scenarios curated to stress-test long-horizon reasoning, making it more challenging than nuScenes for evaluating pure planning performance. In contrast, many two-stage E2E methods on nuPlan report closed-loop metrics that often rely on rule-based post-processing to ensure safety, obscuring their open-loop trajectory prediction capabilities. Our open-loop ADE/FDE on nuPlan provides a cleaner evaluation of the core planning module without such post-processing.
@@
Together, these insights mark a paradigm shift: from reconstructing the world to understanding it. TIWM achieves stronger results not by simulating scenes more faithfully, but by aligning semantics more intelligently. This bridges two long-divided paradigms in embodied AI: world models that emphasize exhaustive synthesis, and VLA systems that are limited to instantaneous reasoning. We show that belief–intent co-evolution, not pixel fidelity, underpins effective planning.
Broader Implications. The “token is all you need” principle transcends autonomous driving. In robotic manipulation, sparse intent tokens could represent dynamic affordances rather than full object geometry. In conversational AI, they could encode evolving discourse states, enabling true foresightful interaction. More broadly, intelligence emerges not from reconstructing environmental details, but from the self-organized balance between what is (belief) and what could be (intent).
Thus, TIWM is not merely a planner—it is a computational model of cognition, where tokens act as atomic units of thought. This work marks a transition from the era of computational intelligence to the era of cognitive intelligence.
VI Conclusion
@@
We challenge the prevailing dogma that high-fidelity future scene reconstruction is indispensable for effective end-to-end autonomous planning. Through extensive experiments on the nuPlan benchmark (720 scenarios, $\sim$11k samples), we demonstrate that a minimalist, cognitively inspired framework—operating on only 16 sparse intent tokens—can achieve state-of-the-art performance. Our Tokenized Intent World Model (TIWM) attains an ADE of 0.382 m without any auxiliary reconstruction supervision, validating that task-driven semantic alignment is not only sufficient but superior to dense modeling paradigms.
@@
This work establishes a new paradigm in autonomous intelligence: planning through understanding. We show that:
  1. 1.
    High-quality perception enables powerful yet simple planners. Reliable BEV inputs allow the system to focus on decision-critical semantics, rendering complex reconstruction losses redundant.
  2. 2.
    Future-aware decoding enables foresightful behavior. Conditioning trajectory generation on autoregressively predicted intent tokens yields a 21.6% performance gain, illustrating that intelligence arises from imagining possible futures, not merely reacting to the present.
  3. 3.
    Explicit reconstruction loss is not only unnecessary—it is detrimental. It introduces gradient conflicts and hinders the emergence of high-level semantic representations, whereas a clean, trajectory-only objective naturally fosters robust and generalizable planning.
Crucially, our model exhibits an emergent cognitive property: temporal fuzziness—the ability to attend flexibly to what matters semantically, rather than aligning rigidly to when it occurs. This, combined with the discovery of cognitive consistency through prolonged training, reveals that intelligence stems not from pixel-level fidelity, but from the coherence of internal cognitive dynamics. The sparse token, in its dual role as belief and intent, acts as a fundamental unit of thought—a dynamic entity undergoing continuous ”quantum state transition” between perceiving the world and imagining its futures.
Beyond autonomous driving, TIWM suggests a universal principle for embodied intelligence: true autonomy arises not from reconstructing the world, but from understanding its semantic essence and simulating its possibilities. By unifying representation, intention, and decision into a single, tokenized loop, our framework bridges the gap between data-driven perception and cognitively inspired reasoning.
Future work will explore closed-loop evaluation, scalability to multi-intent scenarios, and the generalization of belief–intent co-evolution to robotic manipulation and interactive AI. We view TIWM not merely as a planning module, but as a foundational step toward a new generation of agents that do not just react, but truly plan through understanding—ushering in the era of cognitive intelligence.